# LLMs Think, But Not In Your Flow: Reasoning-Level Personalization for Black-Box Large Language Models

### v3

@@ 4.1 Experimental Setup
Datasets. Our framework is evaluated on four peronsalization tasks in NLP including text classification, regression, text generation, and question answering. Three tasks are adopted from the LaMP benchmark [18]— LaMP-2 for movie tagging, LaMP-3 for predicting product ratings, and LaMP-5 for generating scholarly paper titles. Each task is subsampled with 50, 100, and 100 users, respectively, from the original time-based validation splits. Each user’s history is then split into training and test sets in a 9:1 ratio based on chronological order. For personalized question answering, the GlobalOpinionQA (GOQA) dataset [38] is used, where each country is treated as a distinct user group. Following [16], the labels are converted to a single answer by selecting the option with the highest probability, using only instances where this probability exceeds a threshold of 0.8 to ensure label quality. As a result, we obtained 46 user groups (countries) and, for consistency, split each into training and test sets with a 9:1 ratio. Details of datasets are available in Appendix C.1.
Baselines. Our proposed method is compared against a range of baselines for black-box LLM personalization. In addition to the Zero-shot setting, basic personalization methods are considered, including in-context learning (ICL), retrieval-augmented generation (RAG) [18], profile-augmented prompting (PAG) [19], which incorporate user context into input context. HYDRA [17] is included as a plug-and-play personalization method that uses auxiliary modules such as rerankers and adapters to prioritize user-aligned content without modifying the core LLM. Fermi [16], a prompt-based method that iteratively optimizes input prompts using user profiles and feedback from misaligned responses, is also considered. Details of baseline implementations are available in Appendix C.2.
Evaluation metrics. Following [18, 17, 16], accuracy (Acc.) and F1 score (F1) are used for LaMP-2, mean absolute error (MAE) and root mean squared error (RMSE) for LaMP-3, and ROUGE-1 (R-1) [39], ROUGE-L (R-L) [40]for LaMP-5. We report accuracy for the GOQA [38].
Implementation details. For all baselines and our proposed RPM, we adopt GPT-4o-mini [41] as the black-box LLM backbone. We use Contriever [42] as the default retriever for a few-shot example selection, with the number of retrieved examples fixed to 3 in all default settings. During inference, we set the decoding temperature to 0.0 to ensure deterministic outputs. All of the prompt templates used for all tasks are provided in the Appendix G.
Table 1: Overall performance comparison across all benchmarks. For each method, the +CoT variant denotes the use of Chain-of-Thought prompting during inference. RPM  (w/o Reasoning) disables reasoning generation for the target query while keeping retrieved reasoning paths in the input context. Bold and underlined values indicate the best and second-best scores for each metric, respectively.
@@
4.2 Main Results
Table 1 reports the performance for four personalization tasks from the LaMP and GOQA benchmarks. Compared to the Zero-shot, all the baseline methods tend to show improved performance by incorporating user-specific context through few-shot prompting, reranking, or prompt-level optimization. However, these methods focus on the response-level personalization, adapting outputs based on contextual signals without explicitly modeling individual reasoning process.
To examine whether reasoning-inductive prompting can address this limitation, we evaluate the same baselines with chain-of-thought (CoT) [43, 44] reasoning prompts. While CoT is designed to guide the model through intermediate reasoning steps, our results show that it does not consistently improve personalization performance. In several cases, it even leads to degradation, generating longer outputs that are syntactically plausible but misaligned with the user’s actual decision process. These inconsistencies suggest that CoT alone is insufficient: without structured grounding in user-specific reasoning, it merely encourages generic logic rather than meaningful personalization.
In contrast, RPM explicitly targets reasoning-level personalization by constructing user-specific factors from history and using them to generate and retrieve personalized reasoning paths. This allows the model to not only produce outputs aligned with user preferences, but also to reason in ways that reflect the user’s underlying reasoning logic. Across all tasks, RPM consistently outperforms response-level approaches, demonstrating that grounding inference in structured data and personal reasoning is more effective than relying on prompt-level augmentation. We provide additional results in the Appendix D, including backbone variations(Appendix D.5), cost analysis(Appendix D.4), and cross-model transferability(Appendix D.5).
We also evaluate our method without performing explicit reasoning. Notably, some results show that the model can still generate effective responses, suggesting that the provided personalized reasoning examples offer implicit guidance on how to utilize the input, even without explicit reasoning. However, omitting reasoning leads to a consistent performance drop across all tasks, emphasizing the importance of explicit reasoning generation.
Table 2: Ablation study of RPM by adding input components to the zero-shot setting. A simple CoT reasoning example (without using features and factors) is denoted by $r_{q_{i}}^{\text{CoT}}=\mathcal{M}(p_{reason},q_{i};a_{i})$. Bold and underlined values indicate the best and second-best scores for each metric, respectively.
@@ 4.2 Main Results
4.3 Ablation Studies
To evaluate the contribution of each component in our framework to personalization performance, we conduct an ablation study, as summarized in Table 2. Starting from a Zero-shot setting that uses only the target query, we incrementally incorporate user-specific query context: user-specific statistical factors, retrieved query–answer pairs, target query features, and reasoning paths. We observe general trend of performance improvements across tasks at each stage, indicating that both static user representations and contextualized examples contribute meaningfully to personalization. Notably, the inclusion of reasoning paths—even without user-specific personalization—yields substantial gains, demonstrating that reasoning plays a critical role in enhancing personalization performance. The greatest improvement is achieved when personalized reasoning paths are introduced, which model user-specific reasoning logic. These findings validate that effective personalization requires not only incorporating comprehensive user information, but also modeling how individual users reason.
We also observe that even when reasoning paths are constructed solely from raw query-response pairs and used as reasoning examples, the model still achieves strong performance. However, when the reasoning examples and the target query input lack explicit features and factors, performance significantly drops. This suggests that features and factors not only enhance the interpretability of reasoning and facilitate more effective retrieval, but also serve as a crucial foundation and reference point for performing personalized reasoning.
4.4 Evaluating Reasoning-level Interpretability
We evaluate the quality of generated reasoning through a human study conducted on Amazon Mechanical Turk (AMT), comparing RPM against HYDRA and Fermi, both augmented with CoT prompting. We randomly sample 50 examples from each dataset and ask five human annotators per example to assess the reasoning quality across six criteria: Persuasiveness, Reasonability, Faithfulness, Interpretability, Alignment, and Overall Quality. Adapted from prior work [31], these criteria are designed to evaluate both the interpretability of the reasoning and its consistency with user-specific logic. Overall, RPM receives favorable evaluations across all dimensions, with particularly high scores in interpretability and alignment (Figure 3 Left). This suggests that the structured input components employed by RPM are clearly reflected in the model’s outputs through personalized reasoning, thereby enhancing transparency and user-aligned logic. We further evaluate whether the features and factors generated by the LLM fulfill their intended roles. Results indicate that most extracted features are plausibly influential to user responses, and the constructed factors effectively cluster semantically related features in a coherent and meaningful way (Figure 3 Right). Detailed descriptions of each evaluation criterion, the human evaluation protocol, example annotations and case studies are provided in the Appendix C.4.
@@ 4.5 Effectiveness of the Retrieval Strategy
We investigate how different retrieval strategies affect personalization performance, focusing on both the source of examples and the retrieval method in Table 3. Overall, feature-based retrieval from the reasoning-augmented user history, $S_{u}$ yields the most consistent and robust performance across tasks. In the upper section, methods like BM25, Contriever, and the HYDRA reranker retrieve examples based on superficial similarity to the raw query. While they show modest improvements over random selection, their effectiveness is limited—often retrieving topically relevant but decision-irrelevant examples. By contrast, the lower section uses structured retrieval over features and reasoning traces stored in $S_{u}$. These representations better reflect the user’s decision patterns, enabling retrieval that is both contextually and logically aligned. The consistent gains highlight that retrieving from structured, user-specific memory is far more effective than relying on raw queries alone.
@@ 4.6 Impact of User Context Scale on Personalization
Impact of user history length. To assess the effect of user history length on personalization performance, we vary the number of examples from user history used to construct user-specific features and factors. As shown in Figure 4, even a small number of examples enables a meaningful degree of personalization, confirming the feasibility of our approach in low-resource settings. Nevertheless, performance continues to improve with longer histories, indicating that richer context allows for more precise modeling of user decision patterns and reasoning preferences.
@@ 5 Conclusion
We propose RPM, a novel personalization framework for black-box LLMs. RPM is designed to achieve reasoning-level personalization by aligning the model’s generation process with user-specific reasoning logic. Our extensive evaluations across various tasks demonstrate that RPM not only outperforms existing personalization methods, but also that each component is essential for effective personalization. Moreover, the personalized reasoning generated by RPM is interpretable through structured input components and reflects the user’s individual reasoning process by referencing their personal logic. This work represents a paradigm shift in black-box LLM personalization, moving from conventional response-level approaches to a novel reasoning-level personalization framework.
Appendix A Algorithm Details
@@
6:    $\mathcal{G}_{q_{i}}\leftarrow\textsc{FeatureExtract}_{\mathcal{M}}(q_{i})$triangleright extract features with $\mathcal{M}$
@@
8:$\{F^{(m)}\}_{m=1}^{M}\leftarrow\textsc{LLMCluster}_{\mathcal{M}}\!\left(\bigcup_{(q_{i},a_{i})\in\mathcal{H}_{u}}\mathcal{G}_{q_{i}}\right)$triangleright semantic grouping of features with $\mathcal{M}$
9:$\mathcal{C}_{u}\leftarrow\emptyset$triangleright initialize factor set with statistics
@@
12:         $\theta^{(m)}\leftarrow$ComputePropensity($F^{(m)}$) $\triangleright$ compute propensity via Eq. (3)
@@
14:         $\texttt{cov}^{(m)}\leftarrow$ComputeCoverage($F^{(m)}$) $\triangleright$ compute coverage via Eq. (4)
15:         $\texttt{inf}^{(m)}\leftarrow$ComputeInfluence($F^{(m)}$) $\triangleright$ compute influence via Eq. (5) with $\mathcal{M}$
16:         $\texttt{pol}^{(m)}\leftarrow$ComputePolarity($F^{(m)}$) $\triangleright$ compute polarity via Eq. (6) with $\mathcal{M}$
17:         $\theta^{(m)}\leftarrow(\texttt{cov}^{(m)},\texttt{inf}^{(m)},\texttt{pol}^{(m)})$triangleright store three statistics
@@
19:    $\mathcal{C}_{u}\leftarrow\mathcal{C}_{u}\cup\left(F^{(m)},\theta^{(m)}\right)$triangleright append factor + stats to $\mathcal{C}_{u}$
@@
23:    $r_{q_{i}}\leftarrow\textsc{GenerateReasoning}_{\mathcal{M}}(q_{i},\mathcal{G}_{q_{i}},\mathcal{C}_{u},a_{i})$triangleright generate personalized reasoning with $\mathcal{M}$
@@
27:$\mathcal{G}_{q^{\prime}}\leftarrow\textsc{FeatureExtract}_{\mathcal{M}}(q^{\prime})$triangleright extract features from $q^{\prime}$ with $\mathcal{M}$
28:$\mathcal{S}^{\texttt{ret}}_{u,q^{\prime}}\leftarrow$RetrieveTopK($\mathcal{S}_{u},\mathcal{G}_{q^{\prime}}$) $\triangleright$ feature-based retrieval via Eq. (7)
29:$(r_{q^{\prime}},a^{\prime})\leftarrow\textsc{GenerateReasoning}_{\mathcal{M}}(q^{\prime},\mathcal{G}_{q^{\prime}},\mathcal{C}_{u},\mathcal{S}^{\texttt{ret}}_{u,q^{\prime}})$triangleright generate with reasoning-examples with $\mathcal{M}$
@@
Appendix B LLM-based Clustering Details
Among available clustering methods, we adopt the Propose-Assign-Select (PAS) framework [34], which is LLM-based goal-driven explainable clustering method. PAS evaluates both a feature’s name and its context with LLM’s powerful semantic understanding capability, enabling the algorithm to group features that affect the user’s response in similar ways and to provide a natural-language explanation for every resulting group.

### v4

@@ 4.1 Experimental Setup
Datasets. We evaluate our framework on four personalization tasks: text classification, regression, text generation, and question answering. Three tasks are from LaMP (Salemi et al., 2023)—LaMP-2 (movie tagging), LaMP-3 (product rating), and LaMP-5 (paper title generation)—each subsampled with 50, 100, and 100 users, respectively, from the time-based validation splits. User histories are split chronologically into training and test sets (9:1). For personalized QA, we use GlobalOpinionQA (GOQA) (Durmus et al., 2023), treating each country as a user group. Following (Kim & Yang, 2024), we convert labels to a single answer using the highest-probability option, keeping only instances above a 0.8 threshold, yielding 46 user groups. Further details are provided in Appendix D.1.
Baselines. We compare RPM against a range of black-box LLM personalization baselines. In addition to the Zero-shot setting, we include in-context learning (ICL), retrieval-augmented generation (RAG) (Salemi et al., 2023), and profile-augmented prompting (PAG) (Richardson et al., 2023), which incorporate user context into the prompt. HYDRA (Zhuang et al., 2024) is a plug-and-play method that adds rerankers and adapters to prioritize user-aligned content without modifying the core LLM. Fermi (Kim & Yang, 2024) optimizes prompts iteratively using user profiles and feedback from misaligned responses. Details of baseline implementations are available in Appendix D.2.
Evaluation metrics. Following (Salemi et al., 2023; Zhuang et al., 2024; Kim & Yang, 2024), accuracy (Acc.) and F1 score (F1) are used for LaMP-2, mean absolute error (MAE) and root mean squared error (RMSE) for LaMP-3, and ROUGE-1 (R-1) (Lin & Hovy, 2003), ROUGE-L (R-L) (Lin & Och, 2004)for LaMP-5. We report accuracy for the GOQA (Durmus et al., 2023).
Implementation details. All baselines and RPM use GPT-4o-mini (Hurst et al., 2024) as the black-box backbone. Contriever (Izacard et al., 2021) is used for few-shot retrieval with 3 examples by default, and inference is performed with temperature 0.0 for deterministic outputs. For LLM-based clustering (Wang et al., 2023), we use the default parameters without task-specific tuning. More details of implementation details are provided in Appendix D.3
Table 1: Overall performance comparison across all benchmarks. For each method, the +CoT variant denotes the use of Chain-of-Thought prompting during inference. RPM (w/o Reasoning) disables reasoning generation for the target query while keeping retrieved reasoning paths in the input context.
@@
Table 2: Ablation study of RPM by adding input components to the zero-shot setting. A simple CoT reasoning example (without using features and factors) is denoted by $r_{q_{i}}^{\text{CoT}}=\mathcal{M}(p_{reason},q_{i};a_{i})$.
@@
4.2 Main Results
Table 1 reports the performance for four personalization tasks from the LaMP and GOQA benchmarks. Compared to the Zero-shot, all the baseline methods tend to show improved performance by incorporating user-specific context through few-shot prompting, reranking, or prompt-level optimization. However, these methods focus on the response-level personalization, adapting outputs based on contextual signals without explicitly modeling LLM’s reasoning process from user behavior.
To examine whether reasoning-inductive prompting can address this limitation, we evaluate the same baselines with chain-of-thought (CoT) (Wei et al., 2022; Kojima et al., 2022) reasoning prompts. While CoT is designed to guide the model through intermediate reasoning steps, our results show that it does not consistently improve performance. In several cases, it even leads to degradation, generating longer outputs that are syntactically plausible but misaligned with the user’s actual decision. These inconsistencies suggest that CoT alone is insufficient: without structured grounding in reasoning modeled from user behavior, it merely encourages generic logic, not meaningful personalization.
In contrast, RPM explicitly targets reasoning-level personalization by constructing user-specific factors from history and using them to generate and retrieve personalized reasoning paths. This allows the model to not only produce outputs aligned with user preferences, but also to reason in ways that reflect the user’s underlying behavior pattern. Across all tasks, RPM consistently outperforms response-level approaches, demonstrating that grounding inference in structured data and personal reasoning is more effective than relying on prompt-level augmentation.
We also evaluate our method without performing explicit reasoning. Notably, some results show that the model can still generate effective responses, suggesting that the provided personalized reasoning examples offer implicit guidance on how to utilize the input, even without explicit reasoning. However, omitting reasoning leads to a consistent performance drop across all tasks, emphasizing the importance of explicit reasoning generation. Additional experiments demonstrate the framework’s robustness across different backbone models and its strong cross-model transferability (Appendix E.5).
@@ 4.2 Main Results
4.3 Ablation Studies
To evaluate the contribution of each component in our framework to personalization performance, we conduct an ablation study, as summarized in Table 2. Starting from a Zero-shot setting that uses only the target query, we incrementally incorporate user-specific query context: user-specific statistical factors, retrieved query–answer pairs, target query features, and reasoning paths. We observe general trend of performance improvements across tasks at each stage, indicating that both static user representations and contextualized examples contribute meaningfully to personalization. Notably, while incorporating generic reasoning paths provided a performance uplift over simpler variants, this naive approach was still insufficient to outperform the main baseline methods. The greatest improvement is achieved when personalized reasoning paths are introduced, which model user-specific behavior pattern. These findings validate that effective personalization requires not only incorporating comprehensive user information, but also modeling how individual users reason.
We observe that reasoning paths constructed only from raw query–response pairs can still yield strong performance, but performance drops when explicit features and factors are absent. These observations highlight the importance of structured components, which explains why RPM achieves superior personalization. Rather than relying on prompt-level augmentation, RPM employs a systematic framework that transforms raw user history into a structured and quantifiable model of judgment. By extracting response-influential features, aggregating them into user-specific factors, and deriving their statistical significance, the framework provides an interpretable foundation for reasoning. Based on this structured model, RPM generates explicit reasoning paths and employs feature-based retrieval to align inference with user-specific behavior pattern. As confirmed by our ablation study, each stage contributes to improved accuracy, while the framework as a whole ensures robustness and interpretability beyond surface-level personalization. This suggests that features and factors not only enhance the interpretability of reasoning and facilitate more effective retrieval, but also serve as a crucial foundation and reference point for performing personalized reasoning.
4.4 Evaluating Reasoning-level Interpretability
We evaluate the quality of generated reasoning through a human study conducted on Amazon Mechanical Turk (AMT), comparing RPM against HYDRA and Fermi augmented with CoT prompting. We randomly sample 200 examples from datasets and ask five human annotators per example to assess the reasoning quality across six criteria: Persuasiveness, Reasonability, Faithfulness, Interpretability, Alignment, and Overall Quality. These criteria, adapted from (Kim et al., 2025a), evaluate reasoning interpretability and consistency with user-specific behavior logic. Overall, RPM receives favorable evaluations across all dimensions, with particularly high scores in interpretability and alignment (Figure 3 Left). This suggests that the structured input components employed by RPM are clearly reflected in the model’s outputs through personalized reasoning, thereby enhancing transparency and user-aligned behavior pattern. We further evaluate whether the features and factors generated by the LLM fulfill their intended roles. Results indicate that most extracted features are plausibly influential to user responses, and the constructed factors effectively cluster semantically related features coherently and meaningfully (Figure 3 Right). Details of each human evaluation protocol, and case studies are provided in the Appendix D.4 and Appendix H.
@@ 4.5 Effectiveness of the Retrieval Strategy
We investigate how different retrieval strategies affect personalization performance, focusing on both the source of examples and the retrieval method in Table 3. Overall, feature-based retrieval from the reasoning-augmented user history, $S_{u}$ yields the most consistent and robust performance across tasks. In the upper section, methods like BM25, Contriever, and the HYDRA reranker retrieve examples based on superficial similarity to the raw query. While they show modest improvements over random selection, their effectiveness is limited, often retrieving topically relevant but decision-irrelevant examples. By contrast, the lower section uses structured retrieval over features and reasoning traces stored in $S_{u}$. These representations better reflect the user’s choice behavior, enabling retrieval that is both contextually and logically aligned. The consistent gains highlight that retrieving from structured user-specific memory is far more effective than relying on raw queries. Appendix E.3 presents additional experiments on sample-level, feature-level, and factor-guided retrieval.
@@ 4.6 Impact of User Context Scale on Personalization
Impact of user history length. To assess the effect of user history length on personalization performance, we vary the number of examples from user history used to construct user-specific features and factors. As shown in Figure 4, even a small number of examples enables a meaningful degree of personalization, confirming the feasibility of our approach in low-resource settings. Nevertheless, performance continues to improve with longer histories, indicating that richer context allows for more precise modeling of user behaviors and the underlying reasoning structure.
@@
4.7 Cost Analysis
We analyze both inference latency and computational overhead on the GOQA benchmark to address scalability and efficiency. RPM requires only a minor increase in per-user inference time (0.10s vs. 0.04s for ICL/RAG) while delivering substantially higher personalization performance. Our preprocessing cost (\$0.058 per user) and per-instance inference cost (\$0.0037) remain far lower than high-overhead baselines such as Fermi (\$0.32) and HYDRA (\$0.47 + parameter tuning). This shows that RPM offers a practical balance, combining strong personalization and accuracy with modest overhead. For detailed results, including the token usage comparison, please refer to Appendix E.4.
@@ 5 Conclusion
We propose RPM, a novel personalization framework for black-box LLMs. RPM is designed to achieve reasoning-level personalization by aligning the model’s generation process with user-specific decision pattern. Our extensive evaluations across various tasks demonstrate that RPM not only outperforms existing personalization methods, but also that each component is essential for effective personalization. Moreover, the explanations generated by RPM are interpretable, as they are explicitly grounded in structured components derived from the user’s past behavior. This work represents a paradigm shift in black-box LLM personalization, moving from conventional response-level approaches to a novel reasoning-level personalization framework.
Appendix A Notations
Table 4: Notations used throughout the paper.
Notation Description
$u$ A user
$H_{u}=\{(q_{i},a_{i})\}_{i=1}^{N}$ History of user $u$ (queries and responses)
$q,q^{\prime}$ Query and target query
$a,a^{\prime}$ Response and target response
$M$ Black-box LLM used for feature extraction and generation
$G_{q}$ Feature set extracted from query $q$
$F^{(m)}$ A factor obtained by clustering semantically related features
$\theta^{(m)}$ Statistics of factor $F^{(m)}$ (coverage, influence, polarity)
$C_{u}=\{(F^{(m)},\theta^{(m)})\}$ Factor set for user $u$
$r_{q}$ Personalized reasoning path for query $q$
$S_{u}$ Reasoning-augmented memory of user $u$, containing $(q,G_{q},r_{q},a)$
$\text{Coverage}(F^{(m)})$ Fraction of history where factor $F^{(m)}$ appears
$\text{Influence}(F^{(m)})$ Count of cases where $F^{(m)}$ influenced a response
$\text{Polarity}(e,F^{(m)})$ Distribution of polarity labels for features in $F^{(m)}$
Appendix B Algorithm Details
@@
6:  $\mathcal{G}_{q_{i}}\leftarrow\textsc{FeatureExtract}_{\mathcal{M}}(q_{i})$ $\triangleright$ extract features with $\mathcal{M}$
@@
[…]
@@
12:   $\theta^{(m)}\leftarrowComputePropensity$($F^{(m)}$) $\triangleright$ compute propensity via equation 1
@@
14:   $\texttt{cov}^{(m)}\leftarrowComputeCoverage$($F^{(m)}$) $\triangleright$ compute coverage via equation 2
15:   $\texttt{inf}^{(m)}\leftarrowComputeInfluence$($F^{(m)}$) $\triangleright$ compute influence via equation 3 with $\mathcal{M}$
16:   $\texttt{pol}^{(m)}\leftarrowComputePolarity$($F^{(m)}$) $\triangleright$ compute polarity via equation 4 with $\mathcal{M}$
17:   $\theta^{(m)}\leftarrow(\texttt{cov}^{(m)},\texttt{inf}^{(m)},\texttt{pol}^{(m)})$ $\triangleright$ store three statistics
@@
19:  $\mathcal{C}_{u}\leftarrow\mathcal{C}_{u}\cup\left(F^{(m)},\theta^{(m)}\right)$ $\triangleright$ append factor + stats to $\mathcal{C}_{u}$
@@
23:  $r_{q_{i}}\leftarrow\textsc{GenerateReasoning}_{\mathcal{M}}(q_{i},\mathcal{G}_{q_{i}},\mathcal{C}_{u},a_{i})$ $\triangleright$ generate personalized reasoning with $\mathcal{M}$
@@
[…]
@@
Appendix C LLM-based Clustering Details
Among available clustering methods, we adopt the Propose-Assign-Select (PAS) framework (Wang et al., 2023), which is LLM-based goal-driven explainable clustering method. PAS evaluates both a feature’s name and its context with LLM’s powerful semantic understanding capability, enabling the algorithm to group features that affect the user’s response in similar ways and to provide a natural-language explanation for every resulting group.